### v1

# footnotetext: * Equal contribution, Corresponding author

Abstract
Securing personal identity against deepfake attacks is increasingly critical in the digital age, especially for celebrities and political figures whose faces are easily accessible and frequently targeted. Most existing deepfake detection methods focus on general-purpose scenarios and often ignore the valuable prior knowledge of known facial identities, e.g., "VIP individuals" whose authentic facial data are already available. In this paper, we propose VIPGuard, a unified multimodal framework designed to capture fine-grained and comprehensive facial representations of a given identity, compare them against potentially fake or similar-looking faces, and reason over these comparisons to make accurate and explainable predictions. Specifically, our framework consists of three main stages. First, fine-tune a multimodal large language model (MLLM) to learn detailed and structural facial attributes. Second, we perform identity-level discriminative learning to enable the model to distinguish subtle differences between highly similar faces, including real and fake variations. Finally, we introduce user-specific customization, where we model the unique characteristics of the target face identity and perform semantic reasoning via MLLM to enable personalized and explainable deepfake detection. Our framework shows clear advantages over previous detection works, where traditional detectors mainly rely on low-level visual cues and provide no human-understandable explanations, while other MLLM-based models often lack a detailed understanding of specific face identities. To facilitate the evaluation of our method, we built a comprehensive identity-aware benchmark called VIPBench for personalized deepfake detection, involving the latest 7 face-swapping and 7 entire face synthesis techniques for generation. Extensive experiments show that our model outperforms existing methods in both detection and explanation. All code and datasets will be released upon acceptance.
1 Introduction
The rapid advancement of generative AI techniques [46, 54, 63, 68, 47, 12] has led to the widespread creation and dissemination of deepfake content—synthetic media where a person’s identity is manipulated or replaced, often without consent. These techniques severely threaten personal reputation and public trust, especially for high-profile individuals such as celebrities, political figures, and public officials [2]. To this end, protecting one’s identity from such manipulations has become more than just a personal privacy issue—it is a pressing societal concern.
Although deepfake detection has gained attention, most existing methods are designed for general-purpose use [45, 33, 51, 69, 65, 41]. They aim to classify any face image or video as real or fake, without considering whose face is being targeted. In the real world, many scenarios provide access to prior knowledge about the target identity [^1] In this work, we refer to such individuals as ”VIPs” whose identities we aim to protect., such as in the case of public figures or known individuals [15]. This opens a new avenue for detection: Can we leverage the known facial identity to improve both detection and explainability in personalized deepfake detection? Rather than treating all faces the same, identity-aware detection methods [49, 18, 16, 15, 73] focus on the semantic alignment between the input image and the authentic identity. By doing so, they provide more personalized and context-aware predictions, potentially improving both detection performance and explainability.
However, existing identity-aware detectors [15, 73] fail to fully utilize the detailed identity-specific information. They primarily rely on global facial features while neglecting fine-grained semantic details—such as eye shape, facial contours, or other attributes. As shown in Figure 1, an image generated by the latest GPT-4o [43] can appear highly realistic at first glance, but still contains subtle inconsistencies in local facial regions—for example, unusually pronounced eye bags. When the target identity is known and all facial details are available, leveraging these fine-grained discrepancies for detection becomes especially promising.
Figure 1: Comparison between natural images and images generated by GPT-4o [43]. Inconsistencies can be observed in localized facial attributes, such as the eye pouches and facial shapes.
To this end, in this paper, we propose VIPGuard, a unified multimodal framework for detecting and explaining deepfakes targeting specific users. VIPGuard addresses identity-aware deepfake detection by explicitly incorporating known facial identity priors, including both global identity information and detailed structural facial attributes from the VIPs. To achieve this, this paper, for the first time, reformulates forgery detection as a fine-grained face recognition task, targeted at VIP identities. The framework consists of three stages. To leverage both global and local facial information, we use pre-trained face models [57, 13, 10] to extract global facial priors (face similarity scores) and local facial priors (facial attributes). Leveraging these priors, we aim to enable an MLLM to perform forgery detection through semantic comparisons between suspect and authentic faces across visual and textual inputs. We train VIPGuard in three main stages: (1) Fine-tuning an MLLM with a variety of facial attribute data to enhance facial understanding; (2) Performing discriminative learning to distinguish some arbitrary identities from manipulated or similar-looking faces by reasoning; (3) Supporting personalized detection by learning a unique and lightweight VIP token to represent each target identity for customized reasoning.
To enable robust evaluation in personalized deepfake detection, we additionally present VIPBench—a comprehensive and identity-centric benchmark that differs fundamentally from conventional deepfake benchmarks [70, 67]. While existing benchmarks typically treat faces generically and overlook whose identity is being manipulated, VIPBench focuses explicitly on identity-aware scenarios, where prior knowledge of the target individual is available. VIPBench includes 22 specific target identities and a total of 80,080 images, covering both real and forged samples. These forgeries are generated using 14 state-of-the-art methods, spanning 7 face-swapping (FS) and 7 entire-face synthesis (EFS) techniques, providing diverse manipulation types and realistic evaluation settings. By centering evaluation around known identities and incorporating fine-grained annotations, VIPBench allows models to be assessed not only on detection accuracy but also on their ability to leverage identity-specific cues.
Figure 2: Overview of the data collection, VIP-Guard, and Evaluation. VIPGuard’s training and inference pipeline for facial attribute understanding, identity discrimination, and VIP user customization.
Our main contributions are summarized as follows:
  • We introduce a new formulation for personalized deepfake detection that targets specific individuals, casting it as a fine-grained face recognition problem based on both global identity features and detailed facial attributes. This formulation requires only a small number of authentic reference images per VIP user, making it practical for real-world personalized protection scenarios.
  • We propose VIPGuard, a unified multimodal framework for identity-aware deepfake detection and explanation. VIPGuard incorporates pre-trained facial prior models and multimodal large language models, and follows a three-stage pipeline that extracts identity features, performs visual-text reasoning, and supports personalized detection through lightweight identity tokens.
  • We introduce VIPBench, a comprehensive benchmark for evaluating identity-aware deepfake detection. It consists of 22 real-world target identities and 80,080 images generated by 14 state-of-the-art manipulation methods, enabling fine-grained and realistic assessment of personalized detection performance.
2 Related Works
2.1 General Deepfake Detection
Conventional Deepfake Detection
Current deepfake detection faces significant challenges in generalization. To illustrate, researchers have explored a range of approaches, including data augmentation [35, 33, 51, 4], frequency-based cues [45, 40], identity-aware learning [14, 24], disentanglement [71, 69], reconstruction [56], and custom network designs [11]. Data augmentation has proven especially effective for improving generalization—for example, FWA [35] simulates warping artifacts, Face X-ray and SBI [33, 51] target blending boundaries, and SLADD [4] uses adversarial examples to challenge models. Despite these advances, most traditional detectors still offer only binary outputs without human-understandable explanations, leaving users unclear about why a face is classified as fake—limiting trust and explainability.
Deepfake Detection via Multimodal Large Language Model
Vision and language are two core modalities in human perception, driving growing interest in visual-language multimodal learning. In deepfake detection, several studies [28, 50, 72] have explored prompt-based strategies for face forgery analysis, showing that multimodal large language models (MLLMs) offer better explainability than traditional detectors. Others [19, 32, 61, 30, 6, 74] have investigated different MLLMs for explainable detection, while [34] introduced a labeled multimodal dataset to support fine-tuning. However, most of these methods are designed for general-purpose detection and overlook the valuable identity-specific information available in many real-world scenarios.
2.2 Personalized Deepfake Detection for Specific Identity Protection
To detect identity inconsistencies in forged faces, prior works [49, 18, 16, 15, 9] use reference images for personalized deepfake detection. For example, ICT-Ref [15] employs a transformer to detect mismatches between inner- and outer-face regions, while DiffID [73] uses reconstruction-based identity distances to identify fakes. However, these methods mainly rely on global identity features and overlook deeper, user-specific information. This limits their robustness to distribution shifts and prevents them from offering human-understandable explanations.
3 VIPBench: A New Benchmark for Personalized Deepfake Detection
To promote the training and evaluation of personalized deepfake detection, we built a comprehensive identity-aware benchmark called VIPBench for personalized deepfake detection. The training set of VIPBench progressively fine-tunes MLLMs, advancing from basic facial attribute recognition to fine-grained identity inconsistency detection. Moreover, we introduce a new evaluation dataset [^2] We will release the dataset upon acceptance. for identity-aware deepfake detection, a setting that currently lacks sufficient evaluation resources, to assess the effectiveness of different methods. We obtained facial images from LAION-Face [75], CrossFaceID [59], and FaceID-6M [58], subjected to some preprocessing (details in supplementary materials). The main idea of the dataset construction pipeline is described below, while complete details are provided in the supplementary materials
Figure 3: Illustration of the proposed VIPBench, which includes three personalized datasets, (a) Facial Attributes Description Dataset $\mathcal{D}_{FA}$, (b) Identity Discrimination Dataset $\mathcal{D}_{ID}$, and (c) VIPEval $\mathcal{D}_{Eval}$. (d) Some examples of the facial attributes used in the $\mathcal{D}_{FA}$ are also illustrated here, while the full set is available in the supplementary material.
Facial Attributes Description Dataset
We present the Facial Attribute Description Dataset—a multimodal dataset composed of high-resolution facial images paired with rich facial attribute descriptions. The dataset ($\mathcal{D}_{FA}$) is collected about 30,000 high-resolution (more than $1024\times 1024$) images from LAION-Face, to facilitate foundational facial understanding in MLLMs. In addition, an MLLM specialized in facial analysis can act as a captioner to generate descriptive facial attribute information. As illustrated in Figure 3, we obtained detailed attributes (e.g., face shape, skin condition) via MegVii’s official API [^3] https://www.faceplusplus.com.cn/ and subsequently refined them by human experts. Figure 3 (d) provides some examples of these facial attributes, which were then transformed into diverse VQA formats (multiple-choice, short/long answer). The process can be formulated as Eq. 1
$\mathcal{D}_{FA}=\left\{\left(I_{i},\text{VQA}(a_{i1},a_{i2},\ldots,a_{ik})\right)\;\middle|\;(a_{i1},\ldots,a_{ik})\in\mathcal{A}^{k},\;i=1,\ldots,N\right\},$ (1)
where $\mathcal{A}=F(I)$ denotes the set of extracted facial attributes, $F$ is the API, and $I$ is the facial image. The dataset $\mathcal{D}_{FA}$ comprises VQA instances generated from all $k$-tuples of attributes drawn from $\mathcal{A}$. This dataset provides rich supervision for training an MLLM to understand fundamental facial characteristics. The other details about $\mathcal{D}_{FA}$ are shown in the supplementary material.
Identity Discrimination Dataset
By reformulating personalized deepfake detection as a target-face-centric, fine-grained face recognition problem, we constructed the Identity Discrimination Dataset $\mathcal{D}_{ID}$. This dataset, comprising $\mathcal{D}_{ID}^{general}$ and $\mathcal{D}_{ID}^{vip}$, includes facial images and corresponding annotations intended for reasoning about fine-grained identity discrimination. Specifically, $\mathcal{D}_{ID}$ comprises facial image pairs: positive (same-identity, real-real) and negative (all others). As shown in the ‘Data Collection’ part of Figure  2, we built facial pairs centered on VIP users in $D_{ID}^{vip}$, while facial pairs in $\mathcal{D}_{ID}^{{general}}$ were built using arbitrary identities. Given the high resemblance of forged faces to genuine ones, which is challenges for conventional face recognition, we augmented negative samples using SimSwap [5] (face swapping) and Arc2Face [44] (entire face synthesis). These negative pairs include different-identity real-real pairs and real-versus-forgery pairs, with these two categories balanced roughly 1:1. As demonstrated in Figure  3 (b), for each image pair, we first generated a facial attribute list using our fine-tuned Qwen2.5VL-7B model [1] $G$, trained on $\mathcal{D}_{FA}$. Then, we acquired the facial similarity scores by a pre-trained facial recognition model $f$. Gemini 2.5 Pro [^4] Gemini API version in use: 2.5-pro-exp-03-25. [20] was then strictly required to reason identity discrepancies between image pairs based solely on the provided facial attributes $G(I_{i})$ and $G(I_{j})$ and similarity scores $f(I_{i},I_{j})$, ensuring that the analysis was grounded in real features and not influenced by hallucinated content. The above process can be formulated as
$\begin{split}\text{Prompt}&=\text{Gemini}\left(Q,f(I_{i},I_{j}),I_{i},I_{j},G(I_{i}),G(I_{j})\right)\\ \mathcal{D}_{ID}&=\left\{I_{i},I_{j},\text{VQA}\left(I_{i},I_{j},\text{Prompt}\right)\;\middle|\;I_{i},I_{j}\in\mathcal{J}\right\},\end{split}$ (2)
where $I_{i}$ and $I_{j}$ are facial images, $Q$ is the prompt inputted into Gemini, and $\mathcal{J}$ denotes a pre-constructed pool of image-name pairs (see supplemental material for details).
VIPEval: User-Specific Evaluation Dataset
We introduce a user-specific evaluation dataset called VIPEval for assessing user-specific forgery detection performance. Conventional datasets [36, 48] typically lack a sufficient number of high-resolution real images per individual, along with corresponding variations across different face forgery methods, which makes it challenging to evaluate personalized forgery detection approaches. The dataset includes images from rich sources and various resolutions with diverse generation methods, making it more representative of real-world conditions. To address this limitation, we carefully selected 22 unique identities from the previously described image-name pool $\mathcal{J}$, ensuring no overlap with the identities in the $\mathcal{D}_{ID}^{general}$. For each identity, 40-60 real images were collected, 20 reserved for testing in this benchmark $\mathcal{D}_{Eval}$, and the remainder used to construct $\mathcal{D}_{ID}^{vip}$. The benchmark includes reserved real images and a comprehensive set of forged counterparts. For each of the 22 test identities, 420 images are generated per method using seven distinct face-swapping (FS) techniques [5, 52, 27, 62, 64, 60, 17]. Additionally, three open-source entire-face synthesis (EFS) methods [44, 25, 22] generate 10 images per real test image by varying random seeds or prompts, resulting in 200 images per identity. Furthermore, four commercial API-based EFS methods [43, 29, 55, 31] produce 20 images per identity. In total, the dataset comprises 80,080 images.
4 VIPGuard: A Multimodal Framework for Personalized Deepfake Detection
4.1 Overview
We consider a scenario where only a few authentic images of a specific user (e.g., a celebrity, hereafter referred to as a VIP) are available. The objective is to model user-specific facial characteristics to identify suspicious images and protect the VIP’s identity from forgery attacks. To address the challenge, we propose a framework, VIPGuard, which develops an MLLM capable of identifying suspicious images of a specific user, while providing human-understandable explanations based on the user’s unique facial attributes. We reformulate personalized deepfake detection as a fine-grained face recognition problem centered on the protected target, where forgeries are detected through the MLLM’s reasoning over global identity features and detailed facial attributes. To equip the model with this capability, we start from a pre-trained MLLM (Qwen-2.5-VL-7B [1]) and progressively fine-tune it through a three-stage process: Face Attribute Learning, Identity Discrimination, and User-Specific Customization. We describe each stage in detail below.
4.2 Three-Stages Training of VIPGuard
Stage 1: Face Attributes Learning
It is crucial to first enhance the model’s ability to recognize and utilize facial attributes to enable identity protection for VIP users. It is deficient for naive MLLMs to understand human facial features. To this end, we first fine-tune the pre-trained MLLM on $\mathcal{D}_{FA}$ (described in 3). We integrate LoRA [23] modules into the pre-trained MLLM and perform supervised fine-tuning using an autoregressive loss, as defined in Eq. 3:
$\centering L(\bm{\theta})=-\sum_{i=1}^{N}\log\left[p_{\bm{\theta}}(x_{i}\mid x_{i-1},E_{V}(I))\right],\@add@centering$ (3)
where $\bm{\theta}$ denotes the parameters of both the inserted LoRA modules and the original MLLM, $N$ is the length of the output prompt, $x_{i}$ is the $i$-th token to be predicted, $E_{V}$ represents the image encoder of the MLLM, and $I$ is the input image. After training, the MLLM can understand and then recognize facial attributes.
Figure 4: Illustration of the three stages of training the proposed VIPGuard framework.
Stage 2: Identity Discrimination Learning
As illustrated in Figure 1, although the fake face appears realistic at first glance, subtle discrepancies remain in its detailed facial attributes. Therefore, we reformulate personalized deepfake detection as a fine-grained face recognition problem centered on the protected target, where forgeries are detected through the MLLM’s reasoning over global identity features and detailed facial attributes. In this stage, we further fine-tune the MLLM on $\mathcal{D}_{ID}^{{general}}$ for fine-grained facial recognition across face pairs—each consisting of an arbitrary reference face $I{{ref}}$ and an arbitrary query face $I_{{query}}$—thereby equipping it with a foundational ability to identity reason. $\mathcal{D}_{ID}^{\text{general}}$ contains a large number of positive (same-identity) and negative (including both different-identity and forgery) face pairs, each annotated with VQA-style reasoning questions and answers. These annotations support fine-grained identity discrimination by reasoning over both global and local facial prior. Here, we employ face recognition models [57, 13, 10] to provide face similarity scores, serving as a global facial prior, while the local facial prior—i.e., knowledge of facial attributes—has already been incorporated into the fine-tuned MLLM during Stage 1. Specifically, as shown in Figure  4, the reference image $I_{ref}$ and the query image $I_{query}$ are first inputted into the visual encoder $E_{V}$ of our MLLM to obtain the visual token sequences $f_{ref}$ and $f_{query}$, which can be formulated as follows
$\centering f_{ref}=E_{V}(I_{ref}),f_{query}=E_{V}(I_{query}).\@add@centering$ (4)
Pre-trained MLLMs typically lack task-specific optimization for facial recognition, rendering the vanilla visual features $f_{ref}$ and $f_{query}$ suboptimal for facial identity discrimination. Hence, we introduce a Cross-Attention (Cross-Attn) module [37] to capture fine-grained differences in visual tokens between distinct facial images, which is illustrated in Figure 4.
We optimize the MLLM by an autoregressive loss shown as Eq.  5 below.
$\centering L(\bm{{\theta}})=-\sum_{i=1}^{N}\log\left[p_{\bm{{\theta}}}(x_{i}\mid x_{i-1},E_{V}(I_{ref}),E_{V}(I_{query}),g)\right].\@add@centering$ (5)
After training, the model can perform fine-grained facial recognition between any two faces, while also providing detailed explanations of their differences.
Stage 3: User-Specific Customization
Upon completing Stage 2, the MLLM acquires a basic capability for fine-grained facial identity comparison with reasoning. In Stage 3, we further introduce a user-specific customization to enable personalized forgery detection for a given VIP user. To this end, we construct a dataset $\mathcal{D}_{ID}^{{vip}}$ with the same structure as $\mathcal{D}_{ID}^{{general}}$, centered on the VIP’s identity, allowing the model to identify suspicious images by reasoning based on the VIP user’s facial prior. Motivated by the Yo’LLaVA [42], to facilitate lightweight deployment, we incorporate a learnable VIP token trained on $\mathcal{D}_{ID}^{vip}$, denoted as $\bm{\mu}$, which encodes identity-specific features of the VIP user. During this stage, the parameters of the MLLM from Stage 2 are frozen, and only the VIP token is trained on $\mathcal{D}_{ID}^{{vip}}$, thereby refining the model’s ability to perceive and distinguish the target VIP user. As illustrated in Figure 4, the learnable VIP token $\bm{\mu}$, which is a vector of size $32\times d$, is employed to substitute the visual feature representation $f_{ref}$ of the reference image $I_{ref}$. Formally, the prediction procedure can be described as follows
$\centering L(\bm{\mu})=-\sum^{N}_{i}log[p_{\theta}(x_{i}|x_{i-1},\bm{\mu},E_{V}(I_{query}))].\@add@centering$ (6)
Notably, the reference images in $\mathcal{D}_{ID}^{{vip}}$ are not used; only the query image $I_{query}$ and the corresponding reasoning annotation are fed into the MLLM. After training, the MLLM can perform personalized forgery detection for the VIP user by determining whether the identity of query image $I_{{query}}$ belongs to the target individual and providing a user-centric explanation. Furthermore, during inference, we can substitute the corresponding VIP token $\bm{\mu}$ for each user, enabling efficient and accurate detection of malicious facial forgeries without requiring any model retraining. This design supports lightweight and scalable deployment.
5 Experiments
In this section, we present comprehensive experiments to evaluate the effectiveness of our method. For general deepfake detectors, we used models trained on standard deepfake datasets. For ID-aware detectors, we assumed access to 20-30 real images of each VIP user. Detailed experimental settings are provided in the supplementary material.
Evaluation Metrics
For evaluation, we report three standard metrics: Area Under the Curve (AUC), Equal Error Rate (EER), and Accuracy (ACC). AUC measures the model’s ability to distinguish between positive and negative classes across all thresholds. EER represents the point where the false acceptance rate is equal to the false rejection rate. ACC denotes the proportion of correct predictions. We use AUC and EER to compare our method following prior work [67, 70]. Moreover, ACC is used to compare with API-based commercial generators (e.g., GPT-4o [^5] GPT4o API version in use: GPT-4o-2024-08-06 [43]).
5.1 Evaluations on the proposed VIPEval and Other Existing Benchmarks
Comparison with Deepfake Detection Methods on VIPEval
We evaluated general deepfake detectors, ID-aware detectors, and our proposed method on VIPEval, reporting results for face swapping (Table  1) detection and entire faces synthesis detection(Table  2). Our method significantly improves the detection of all facial forgery types. General deepfake detectors (e.g., Effort [66]) excel at face swapping detection but fail to detect more realistic forgeries from commercial APIs due to unseen artifacts. ID-aware detectors, unlike general detectors reliant on low-level cues, leverage identity-related semantic consistency for robust detection across diverse generation techniques and data. However, existing ID-aware methods (e.g., DiffID, ICT-Ref) struggle with fully synthesized faces due to their reliance on global facial features, a limitation amplified by advancing forgery techniques. Conversely, VIP-Guard captures identity-specific cues by jointly leveraging global facial representations and local attribute analysis, enabling robust protection of VIP users with few real photos across diverse forgery methods.
Table 1: Evaluation of generalization performance (AUC (%) / EER (%)) for face swapping detection on VIPEval.
Methods BlendFace [52] Ghost3 [17] HifiFace [60] InSwap [27] MobileSwap [64] SimSwap [5] UniFace [62]
Xception [8] 53.89 / 46.59 61.08 / 42.27 71.70 / 34.09 64.79 / 38.64 98.12 / 7.50 64.91 / 38.41 59.91 / 42.50
EfficientNet [53] 33.94 / 60.91 43.52 / 54.32 61.22 / 42.95 37.34 / 58.64 89.67 / 64.45 56.55 / 45.00 46.53 / 50.23
UCF [69] 64.31 / 38.18 65.92 / 35.68 65.62 / 36.82 66.02 / 34.32 87.33 / 20.00 63.11 / 36.82 67.25 / 34.32
ProDet [7] 59.84 / 42.95 53.18 / 47.73 56.48 / 46.36 35.91 / 59.32 73.01 / 34.32 56.27 / 45.68 49.34 / 50.00
RECCE [3] 56.11 / 45.00 61.53 / 39.54 56.93 / 45.91 58.85 / 43.41 88.40 / 19.32 60.05 / 42.95 63.01 / 39.10
CDFA [39] 59.86 / 45.23 70.50 / 36.82 85.43 / 24.09 70.92 / 36.59 98.75 / 5.00 71.22 / 35.91 73.30 / 34.77
RepDFD [38] 70.34 / 35.00 78.76 / 28.18 80.09 / 27.50 71.66 / 34.32 95.76 / 10.45 79.40 / 27.95 82.36 / 25.45
Effort [66] 91.87 / 17.61 95.28 / 13.07 96.60 / 10.23 85.53 / 23.86 96.23 / 13.07 97.16 / 8.52 92.48 / 13.64
DiffID [73] 83.33 / 24.47 66.02 / 38.38 87.23 / 20.33 66.22 / 38.06 75.71 / 30.89 72.03 / 33.91 83.30 / 24.40
ICT-Ref [15] 88.67 / 13.37 86.52 / 15.51 85.90 / 14.05 84.34 / 17.87 87.45 / 13.52 86.57 / 14.65 82.73 / 19.04
VIPGuard 99.48 / 0.81 97.97 / 4.39 99.63 / 0.61 96.40 / 8.24 99.55 / 1.01 99.43 / 1.02 99.69 / 0.26
Table 2: Evaluation of generalization performance (AUC (%) / EER (%)) for entire face synthesis detection on VIPEval.
Method Open-Source Commercial-API
ConsistentID [25] Arc2Face [44] PuLID [22] GPT-4o [43] Jimeng AI [29] TongYi [55] Kling AI [31]
Xception 42.02 / 54.77 51.87 / 48.86 59.23 / 44.09 58.13 / 46.36 57.77 / 44.55 34.36 / 62.05 44.34 / 54.32
EfficientNet 33.81 / 61.14 44.96 / 50.91 47.19 / 50.23 75.35 / 28.86 55.04 / 45.00 41.15 / 55.68 42.40 / 54.32
UCF [69] 62.16 / 40.91 56.62 / 46.36 54.16 / 46.36 59.06 / 42.73 71.08 / 32.78 82.38 / 22.73 63.31 / 40.23
ProDet [7] 63.68 / 37.95 59.62 / 40.91 67.02 / 36.82 59.23 / 40.91 59.71 / 42.73 89.80 / 18.41 72.53 / 32.73
RECCE [3] 68.56 / 36.82 57.79 / 47.72 63.85 / 40.00 83.63 / 22.73 69.00 / 35.23 97.00 / 7.50 70.94 / 37.72
CDFA [39] 77.62 / 30.00 67.09 / 39.09 67.93 / 37.50 73.46 / 32.73 71.98 / 34.09 90.32 / 17.95 77.47 / 30.00
RepDFD [38] 83.52 / 24.32 61.67 / 41.59 74.65 / 32.27 73.62 / 32.27 62.78 / 40.91 93.80 / 14.09 60.10 / 43.41
Effort [66] 58.68 / 47.35 57.03 / 44.89 56.31 / 44.89 49.63 / 48.30 63.60 / 40.34 82.93 / 23.86 56.84 / 44.89
DiffID [73] 75.85 / 32.01 78.47 / 28.86 70.36 / 35.72 45.26 / 52.97 64.29 / 39.35 84.66 / 23.58 69.51 / 35.56
ICT-Ref [15] 63.15 / 39.84 70.27 / 32.84 72.36 / 31.93 58.59 / 40.93 65.36 / 35.73 74.88 / 24.41 50.05 / 45.98
VIPGuard 99.69 / 0.45 98.05 / 4.80 98.96 / 1.90 89.03 / 16.14 97.04 / 5.36 99.76 / 0.23 99.27 / 1.25
Table 3: Comparison (ACC (%)) of our method and other MLLMs on the VIPEval.
Method BlendFace HifiFace MobileSwap UniFace ConsistentID Arc2Face
GPT-4o-2024-08-06[43] 71.85 84.36 94.35 70.72 46.70 50.90
Gemini-2.5-pro-exp-03-25 [20] 80.87 82.68 91.75 83.58 69.85 79.98
Qwen2.5VL 7B [1] 50.05 49.86 49.99 50.02 49.51 49.69
LLaMA3.2Vision 11B [21] 44.76 43.93 45.71 54.09 38.56 43.61
FFAA [26] 88.60 89.95 91.36 89.30 63.53 59.89
VIPGuard 95.51 95.82 96.71 95.88 95.91 89.71
Comparison with Other LLM-based Methods on VIPEval
This experiment evaluated the detection capabilities of Multimodal Large Language Models (MLLMs) on VIPEval by comparing various MLLM-based methods, including FFAA [26], an MLLM specialized for face forgery detection, alongside naive MLLMs. Detection performance was measured using Accuracy (ACC) due to the API’s binary (Real/Fake) output. As presented in Table  3, our method consistently outperforms other approaches across all forged image types, demonstrating the effectiveness of identity-specific semantic detection.
Table 4: Evaluation of frame-level performance (AUC (%) / EER (%)) in CelebDF [36] and DF40 [67] under one-shot setting. For each identity, only a single real image is available.
Method CelebDF DF40
BlendFace E4S FaceDancer FSGAN SimSwap UniFace InSwap
Diff-ID 86.66 / 21.79 76.83 / 28.99 84.79 / 24.07 81.04 / 27.02 82.88 / 23.77 64.87 / 38.28 84.04 / 22.53 59.62 / 43.19
ICT-Ref 81.86 / 26.18 76.47 / 30.63 83.33 / 24.93 91.57 / 16.38 73.62 / 32.95 82.41 / 25.62 80.34 / 27.07 69.91 / 35.98
VIPGuard 87.96 / 20.84 81.87 / 26.23 89.74 / 19.63 86.12 / 23.61 86.44 / 20.87 83.65 / 24.92 86.70 / 22.57 70.97 / 34.15
Figure 5: Visual illustration of the analysis of VIP-Guard detecting anomalous local facial attributes for EFS (left) and FS (right).
One-shot Performance on other Deepfake detection datasets
This experiment evaluates our method against several existing ID-aware approaches on established benchmark datasets. The evaluation specifically employed the challenging CelebDF [36] dataset and a CelebDF-related subset of DF40 [67]. Owing to limited real sample diversity in these datasets, our method was evaluated in a one-shot setting, using a single real image per user. To mitigate similarity from shared video sources, reference and test images were sampled from different videos. As shown in Table 4, our method still achieves competitive performance compared to other approaches. Notably, due to limited real image availability, these results were obtained using VIP-Guard at Stage 2, omitting Stage 3. The results also verify the effectiveness of Stage 2 of our method.
Visual Examples of Our Model Explanations
Figure 5 presents two representative examples illustrating VIPGuard’s ability to detect anomalous local facial attributes in deepfake images generated by entire face synthesis (EFS, left) and face swapping (FS, right) techniques. Despite the high global facial similarity between the real and fake images, VIPGuard accurately identifies subtle but critical differences in localized regions such as the eyes, lips, and skin textures. For instance, discrepancies in eye size, lip thickness, and specific skin features (e.g., glabella wrinkles, eye pouches, moles, and crow’s feet) are effectively captured by our model. These examples highlight two key strengths of VIPGuard: (1) its ability to detect forgery-induced anomalies in local facial regions where current generation methods often fail, and (2) its capacity to reason over these attributes for reliable identity verification, even when global facial appearance is highly similar.
6 Conclusion
This paper proposes VIPGuard, addressing a critical gap in deepfake detection by leveraging known facial identities to enable personalized, accurate, and explainable detection. Unlike traditional detectors that mainly rely on low-level visual artifacts or general-purpose MLLMs lacking identity awareness, VIPGuard integrates fine-grained attribute learning, identity-level discriminative training, and user-specific customization through a unified multimodal framework. Combined with our newly proposed VIPBench, which enables rigorous evaluation across diverse and advanced forgery types, our approach demonstrates clear superiority in both detection performance and explainability, offering a robust and scalable solution for safeguarding high-risk individuals against identity-based deepfakes.
Content Structure of the Appendix.
Due to page constraints, we include additional analyses and experiments in the Appendix, containing comprehensive ablation studies, more visual examples of model explanations, details of dataset construction, an ethical statement, and limitations and future work. For further details, please refer to the Appendix.